# Live-SWE-agent: Can Software Engineering Agents Self-Evolve on the Fly?

### v1

@@ Abstract
Large Language Models (LLMs) are reshaping almost all industries, including software engineering. In recent years, a number of LLM agents have been proposed to solve real-world software problems. Such software agents are typically equipped with a suite of coding tools and can autonomously decide the next actions to form complete trajectories to solve end-to-end software tasks. While promising, they typically require dedicated design and may still be suboptimal, since it can be extremely challenging and costly to exhaust the entire agent scaffold design space. Recognizing that software agents are inherently software themselves that can be further refined/modified, researchers have proposed a number of self-improving software agents recently, including the Darwin-Gödel Machine (DGM). Meanwhile, such self-improving agents require costly offline training on specific benchmarks and may not generalize well across different LLMs or benchmarks. In this paper, we propose Live-SWE-agent, the first live software agent that can autonomously and continuously evolve itself on-the-fly during runtime when solving real-world software problems. More specifically, Live-SWE-agent starts with the most basic agent scaffold with only access to bash tools (e.g., mini-SWE-agent), and autonomously evolves its own scaffold implementation while solving real-world software problems. Our evaluation on the widely studied SWE-bench Verified benchmark shows that Live-SWE-agent can achieve an impressive solve rate of 75.4% without test-time scaling, outperforming all existing open-source software agents and approaching the performance of the best proprietary solution. Moreover, Live-SWE-agent outperforms state-of-the-art manually crafted software agents on the recent SWE-Bench Pro benchmark, achieving the best-known solve rate of 45.8%. More details at: [Uncaptioned image] https://github.com/OpenAutoCoder/live-swe-agent
@@
Figure 1: Live-SWE-agent on SWE-bench Verified (sbv) and SWE-Bench Pro (deng2025swe).
@@ 1 Introduction
We have evaluated Live-SWE-agent on the widely used SWE-bench Verified benchmark sbv and the more challenging SWE-Bench Pro deng2025swe. Without any test‑time scaling, Live-SWE-agent achieves 75.4% resolve rate on Verified and 45.8% on Pro, surpassing state-of-the-art open‑source baselines and approaching the best commercial agent systems. Our ablations and tooling analyses reveal that (1) custom tool creation materially improves effectiveness (higher solve rates) with minimal overhead, (2) benefits persist across different state-of-the-art LLM backends, with better results on stronger models, demonstrating its promising future as LLMs’ capabilities rapidly evolve, and (3) synthesized tools include both generic utilities and task-aligned specializations that benefit issue-specific problem-solving.
@@
  • State-of-the-art open-source agent. On SWE-bench Verified and SWE-Bench Pro, Live-SWE-agent reaches 75.4% and 45.8% solve rates respectively, outperforming leading open-source agents and approaching commercial systems, without any test-time scaling. To our knowledge, our SWE-Bench Pro result is the best reported to date.
@@ 3 Experimental Setup
Implementation. While Live-SWE-agent is general across different agent scaffolds, we implement Live-SWE-agent on top of the popular mini-SWE-agent (yang2024sweagent) framework. As such, we retain the hyperparameters used in mini-SWE-agent by default (i.e., temperature 0.0, maximum step limit of 250, and maximum cost of \$3 per issue). We choose mini-SWE-agent to build on since it is not only simplistic (with $\sim$100 lines of code and only accessing bash commands) but also widely used. By default, we use Claude 4.5 Sonnet (claude-sonnet-4-5-20250929(sonnet4.5) in our experiments and sample one patch per issue.
@@ 4.1 Main Results
We also demonstrate the comparison between Live-SWE-agent and the best-performing agentic tools, including both state-of-the-art open-source solutions and proprietary commercial scaffolds, on SWE-bench Verified (Figure 1). We report both the base agent performance and, when applicable, the result with test-time scaling. We observe that Live-SWE-agent using Claude 4.5 Sonnet achieves a solve rate of 75.4% without test-time scaling, outperforming all open-source agents on the SWE-bench Verified leaderboard [^1] https://www.swebench.com at the time of writing and achieving performance comparable to state-of-the-art commercial solutions.
@@ 4.4 Discussion and Future Work
Impact on LLM evaluation. Our experimental results on multiple benchmarks using multiple LLMs demonstrate that Live-SWE-agent can achieve state-of-the-art performance compared with open-source solutions and comparable results with proprietary tools. This shows that Live-SWE-agent can provide an effective scaffold for unified evaluation of LLM performance on solving software development issues. Instead of using complex or proprietary agent scaffolds, Live-SWE-agent provides a simplistic and lightweight approach that can be easily added on top of any agent design to evaluate LLMs. Moreover, Live-SWE-agent not only evaluates the issue solving ability, but it can also test the tool-creation capability of LLMs and agents. Tools are one of the most important aspect of an agent and directly impacts the issue solving performance. The ability for LLMs to create custom tools, especially for solving complex software development issues, have not been extensively evaluated. To this end, we have already seen some interesting results in our evaluation where weaker LLMs lack high-level reasoning to create useful tools on the fly (see Section 4.3). Live-SWE-agent provides a unique framework to jointly evaluate both the tool creation and issue resolution ability of LLMs.
@@ 6 Conclusion
In this paper, we have proposed Live-SWE-agent, the first live software agent that can autonomously and continuously evolve itself on-the-fly during runtime when solving real-world software problems. More specifically, Live-SWE-agent starts with the most basic agent scaffold with only access to bash tools, and autonomously evolves its own scaffold implementation while solving real-world software problems. Our evaluation on the widely studied benchmarks (such as SWE-bench Verified and SWE-bench Pro) demonstrated that Live-SWE-agent can approach or even outperform state-of-the-art manually design software agents, demonstrating a promising future for live and self-evolving software agents.
@@ Appendix A Additional experimental settings
We now describe additional experimental settings used by Live-SWE-agent for evaluation. As discussed in Section 3, we implement Live-SWE-agent on top of the mini-SWE-agent framework. We use the same default hyperparameter of mini-SWE-agent at the time of writing (i.e., temperature 0.0, maximum step limit of 250, and maximum cost of \$3 per issue). For any OpenAI models in our experiments (e.g., GPT-5, GPT-5-Mini, GPT-5-Nano), since temperature 0.0 sampling is not supported, we use temperature of 1. To the best of our knowledge, this is also the case used in mini-SWE-agent when they use OpenAI models.

### v2

@@ Abstract
Large Language Models (LLMs) are reshaping almost all industries, including software engineering. In recent years, a number of LLM agents have been proposed to solve real-world software problems. Such software agents are typically equipped with a suite of coding tools and can autonomously decide the next actions to form complete trajectories to solve end-to-end software tasks. While promising, they typically require dedicated design and may still be suboptimal, since it can be extremely challenging and costly to exhaust the entire agent scaffold design space. Recognizing that software agents are inherently software themselves that can be further refined/modified, researchers have proposed a number of self-improving software agents recently, including the Darwin-Gödel Machine (DGM). Meanwhile, such self-improving agents require costly offline training on specific benchmarks and may not generalize well across different LLMs or benchmarks. In this paper, we propose Live-SWE-agent, the first live software agent that can autonomously and continuously evolve itself on-the-fly during runtime when solving real-world software problems. More specifically, Live-SWE-agent starts with the most basic agent scaffold with only access to bash tools (e.g., mini-SWE-agent), and autonomously evolves its own scaffold implementation while solving real-world software problems. Our evaluation on the widely studied SWE-bench Verified benchmark shows that Live-SWE-agent can achieve an impressive solve rate of 77.4% without test-time scaling, outperforming all existing software agents, including the best proprietary solution. Moreover, Live-SWE-agent outperforms state-of-the-art manually crafted software agents on the recent SWE-Bench Pro benchmark, achieving the best-known solve rate of 45.8%. More details at: [Uncaptioned image] https://github.com/OpenAutoCoder/live-swe-agent
@@
Figure 1: SWE-bench Verified and SWE-Bench Pro single attempt (w/o test-time scaling) results.
@@ 1 Introduction
We have evaluated Live-SWE-agent on the widely used SWE-bench Verified benchmark sbv and the more challenging SWE-Bench Pro deng2025swe. Without any test‑time scaling, Live-SWE-agent achieves 77.4% resolve rate on Verified and 45.8% on Pro, surpassing state-of-the-art open‑source baselines and even the best commercial agent systems. Our ablations and tooling analyses reveal that (1) custom tool creation materially improves effectiveness (higher solve rates) with minimal overhead, (2) benefits persist across different state-of-the-art LLM backends, with better results on stronger models, demonstrating its promising future as LLMs’ capabilities rapidly evolve, and (3) synthesized tools include both generic utilities and task-aligned specializations that benefit issue-specific problem-solving.
@@
  • State-of-the-art open-source agent. On SWE-bench Verified and SWE-Bench Pro, Live-SWE-agent reaches 77.4% and 45.8% solve rates respectively, outperforming leading open-source agents and commercial systems, without any test-time scaling. To our knowledge, our SWE-Bench Pro result is the best reported to date.
@@ 3 Experimental Setup
Implementation. While Live-SWE-agent is general across different agent scaffolds, we implement Live-SWE-agent on top of the popular mini-SWE-agent (yang2024sweagent) framework. As such, we retain the hyperparameters used in mini-SWE-agent by default (i.e., temperature 0.0, maximum step limit of 250, and maximum cost of \$3 per issue). We choose mini-SWE-agent to build on since it is not only simplistic (with $\sim$100 lines of code and only accessing bash commands) but also widely used. Unless otherwise stated, we use Claude 4.5 Sonnet (claude-sonnet-4-5-20250929(sonnet4.5) in our experiments and sample one patch per issue.
@@ 4.1 Main Results
We also demonstrate the comparison between Live-SWE-agent and the best-performing agentic tools, including both state-of-the-art open-source solutions and proprietary commercial scaffolds, on SWE-bench Verified (Figure 1). We report both the base agent performance and, when applicable, the result with test-time scaling. We observe that Live-SWE-agent using Gemini 3 Pro achieves a solve rate of 77.4% without test-time scaling, outperforming all existing agents on the SWE-bench Verified leaderboard [^1] https://www.swebench.com including state-of-the-art commercial solutions at the time of writing.
@@ 4.4 Discussion and Future Work
Impact on LLM evaluation. Our experimental results on multiple benchmarks using multiple LLMs demonstrate that Live-SWE-agent can achieve state-of-the-art performance. This shows that Live-SWE-agent can provide an effective scaffold for unified evaluation of LLM performance on solving software development issues. Instead of using complex or proprietary agent scaffolds, Live-SWE-agent provides a simplistic and lightweight approach that can be easily added on top of any agent design to evaluate LLMs. Moreover, Live-SWE-agent not only evaluates the issue solving ability, but it can also test the tool-creation capability of LLMs and agents. Tools are one of the most important aspect of an agent and directly impacts the issue solving performance. The ability for LLMs to create custom tools, especially for solving complex software development issues, have not been extensively evaluated. To this end, we have already seen some interesting results in our evaluation where weaker LLMs lack high-level reasoning to create useful tools on the fly (see Section 4.3). Live-SWE-agent provides a unique framework to jointly evaluate both the tool creation and issue resolution ability of LLMs.
@@ 6 Conclusion
In this paper, we have proposed Live-SWE-agent, the first live software agent that can autonomously and continuously evolve itself on-the-fly during runtime when solving real-world software problems. More specifically, Live-SWE-agent starts with the most basic agent scaffold with only access to bash tools, and autonomously evolves its own scaffold implementation while solving real-world software problems. Our evaluation on the widely studied benchmarks (such as SWE-bench Verified and SWE-bench Pro) demonstrated that Live-SWE-agent outperforms state-of-the-art manually design software agents, demonstrating a promising future for live and self-evolving software agents.
@@ Appendix A Additional experimental settings
We now describe additional experimental settings used by Live-SWE-agent for evaluation. As discussed in Section 3, we implement Live-SWE-agent on top of the mini-SWE-agent framework. We use the same default hyperparameter of mini-SWE-agent at the time of writing (i.e., temperature 0.0, maximum step limit of 250, and maximum cost of \$3 per issue).
For Gemini 3 Pro, we use temperature of 1 since it is recommended by the developers [^4] https://ai.google.dev/gemini-api/docs/gemini-3. For any OpenAI models in our experiments (e.g., GPT-5, GPT-5-Mini, GPT-5-Nano), since temperature 0.0 sampling is not supported, we use temperature of 1. To the best of our knowledge, this is also the case used in mini-SWE-agent when they use OpenAI models.